OpenStudy (anonymous):

What is the surface area of a cone with a radius of 12 centimeters and a slant height of 8 centimeters? use 3.14 for pie. a. 502.4cm2 b. 753.6cm2 c. 803.84cm2 d. 1205.76cm2

OpenStudy (anonymous):

a b c or d

OpenStudy (anonymous):

sorry i can't just give you the answer.. the formula is in the link i posted above. all you have to do is to plug in what you have into the formula and play it on a calculator.

OpenStudy (anonymous):

it wont let me go

OpenStudy (anonymous):

\[\Large A=\pi rs +\pi r^2\] A=Surface area of the cone r=radius of the cone s=the slant height of the cone